# M3D: Manifold-based Domain Adaptation with Dynamic Distribution for Non-Deep Transfer Learning in Cross-subject and Cross-session EEG-based Emotion Recognition

### v2

@@ IV-D Experimental Results on Cross-Subject Cross-Session
IV-E Experimental Results on clinical EEG database
To assess the generalizability of the proposed M3D model in real-world clinical applications, we evaluate its performance on the MDD database (n=56). Following previous work [77], we adopt the cross-subject single-session 10-fold cross-validation protocol for fair comparison. The validation results are presented in Table VII, demonstrating consistent performance across various initial classifiers. In particular, the highest accuracy is achieved using DT as the initial classifier, reaching 82.16% in the EC session and 78.49% in the EO session. For comparative analysis, we report performance differences against existing methods in Table VIII. The proposed M3D model outperforms the best-performing traditional machine learning approach by 7.95% and 4.28% in the EC and EO sessions, respectively. Among deep learning methods, DCORAL [75] achieves the highest reported accuracy in prior studies, with 70.88±10.53% in the EC session and 74.49±08.17% in the EO session. M3D surpasses these results by 11.28% and 4% in the EC and EO sessions, respectively, demonstrating its superior ability to capture the complex patterns associated with MDD. These results underscore the effectiveness of M3D in real-world clinical settings, establishing it as a promising tool for practical clinical applications.
TABLE VII: Cross-subject single-session 10 fold cross-validation on MDD with different initial classifiers.
@@ IV-E Experimental Results on clinical EEG database
EO Accuracy 74.92 74.92 78.49 74.92 74.92 74.92
@@
TABLE VIII: Performance comparison on MDD using cross-subject 10-fold cross-validation. Models reproduced by the authors are indicated with an asterisk (*), while the second-best traditional machine learning method’s results are underlined. EC: closed-eye; EO: open-eye.
@@
      M3D       82.16/11.35       78.49/13.79
TABLE IX: Performance comparison without classifier learning and ensemble learning, with only classifier learning, and with both classifier learning and ensemble learning.
Manifold Feature Transformation Dynamic Distribution Alignment and Classifier Learning Ensemble Learning Accuracy F1-Score 48.27/14.39 35.42 69.51/16.40 56.53 71.01/13.42 59.41 79.66/09.34 69.94 84.57/09.49 76.11
V Discussion
V-A Ablation Study
@@
Table IX presents the results of the ablation study, demonstrating that the proposed M3D model achieves the best performance by effectively integrating all modules. (1) Manifold Feature Transformation Only: When only the manifold feature transformation module is used, the model performance is the lowest, with a classification accuracy of 44.73±14.39%. (2) Dynamic Distribution Alignment and Classifier Learning Only: Using only dynamic distribution alignment and classifier learning yields better results than using only manifold feature transformation, achieving an accuracy of 69.51±16.40%. (3) Without Manifold Feature Transformation: Adding ensemble learning to dynamic distribution alignment and classifier learning further improves accuracy to 71.01±13.42%, reflecting a 1.5% improvement. This highlights the effectiveness of ensemble learning in enhancing EEG-based emotion recognition performance. However, compared to the full model, removing the manifold feature transformation module results in a significant accuracy drop from 84.57±09.49% to 71.01±13.42%. This underscores the importance of feature transformation before distribution alignment in effectively reducing domain differences and improving model performance. (4) Without Ensemble Learning: Removing the ensemble learning module negatively impacts the model, leading to a reduced accuracy of 79.66±09.34%. Additionally, features in the Grassmann manifold space prove more robust than those in the original feature space, improving distribution alignment. All four ablation models differ significantly from the full model (p $<$ 0.0001, paired t-test, FDR corrected).
TABLE X: Performance comparison using TCA and PCA.
         Methods          PCA          TCA
         Accuracy          52.93          84.57
         F1-Score          50.55          76.11
TABLE XI: Performance comparison with different $\mu$ values.

### v3

@@ 4.2 Experimental Protocols
Table 8: Performance comparison on SEED-V using cross-subject cross-session LOSO CV. The second-best traditional machine learning method result is underlined. Models reproduced by the authors are indicated with an asterisk (*).
Methods $P_{acc}$ Methods $P_{acc}$
Traditional machine learning methods
TCA*[18] 34.15/08.04 SA*[23] 41.20/11.95
GFK*[25] 35.94/09.51 SVM*[38] 46.72/12.50
Adaboost*[40] 47.46/12.82 KNN*[51] 35.00/09.38
CORAL*[53] 46.23/12.58 KPCA*[52] 32.23/12.09
Deep learning methods
DGCNN[57] 37.38/08.64 DANN [57] 46.87/10.09
ADDA[57] 47.97/09.43 GMSS[57] 46.34/09.65
MSLTE[57] 50.12/08.59 A-LSTM[58] 40.34/08.68
BDGLS[58] 59.59/04.75 IAG[58] 59.68/09.44
ECLGCNN[58] 61.58/10.35 DAN*[14] 59.32/10.30
DCORAL*[59] 57.41/09.65 DDC*[60] 59.44/07.04
M3D 59.15/09.30
4.3 Experimental Results on Cross-Subject Single-Session
The experimental results on SEED, SEED-IV and SEED-V database under the cross-subject single-session cross-validation are reported in Table 1. For comprehensive evaluation, we compare our method with existing approaches under identical experimental conditions. The comparison methods included in each table are selected based on the availability of reported results under the corresponding dataset and evaluation protocol to ensure fair and consistent evaluation. The baseline methods selected for comparison include a range of classical and recent approaches, covering representative state-of-the-art techniques published in the field of EEG-based domain adaptation. Results marked with * are independently reproduced by the authors under standardized preprocessing and evaluation protocols to ensure comparability.
For the SEED database, the best performance is obtained when DT is adopted as the initial classifier, where the corresponding accuracy, specificity and precision are 84.57%, 88.78% and 76.94%. We compare the obtained experimental results with the existing literature, as shown in Table 2. The comparison results show that the proposed non-deep transfer learning model (M3D) demonstrates superior performance. In comparison to the best results obtained by the traditional machine learning methods, our model shows a performance enhancement of 8.26%. Furthermore, it performs comparably or even surpasses some deep learning methods, such as ADA [16], TANN [47], P-GCNN [48] and DG-DANN [43].
For the SEED-IV database, the best accuracy is obtained when DT is used as the initial classifier, where the corresponding accuracy, sensitivity, F1-score, AUROC and NPV are 60.94%, 78.10%, 50.49%, 73.68% and 88.02%, respectively. The performance comparison results with the existing methods are reported in Table 3. Compared to the best results reported in the traditional machine learning methods, the proposed M3D demonstrates a slight performance advantage of 0.72%. Furthermore, it performs comparably or even surpasses some deep learning methods, including ScalingNet [46]and MS-MDA [50].
For the SEED-V database, the best accuracy is obtained when DT is used as the initial classifier, where the corresponding accuracy, specificity, precision, F1-score and AUROC are 65.25%, 66.16%, 16.6%, 25.76% and 78.46%. As shown in Table 4, the M3D model outperforms the best-performing traditional machine learning baseline by a notable 14.15%, clearly demonstrating its superior classification capability.
4.4 Experimental Results on Cross-Subject Cross-Session
To evaluate the efficiency and stability of the proposed M3D in managing variations in EEG data collected from the same subject at different times, we also conduct a model validation under the cross-subject cross-session leave-one-subject-out cross-validation protocol. The experimental results are reported in Table 5, when different initial classifiers are used. The best model performance of SEED database is achieved when DT is adopted as the initial classifier, where the accuracy is 77.05%, and the other six indicators, except for NPV, are also the highest. The performance results against the existing methods are compared in Table 6. It shows, the proposed model demonstrates an accuracy improvement of 5.18% in comparison to traditional machine learning methods. Further, it exhibits performance that is comparable, to some extent, with deep learning methods, showing the effectiveness of the proposed M3D in managing variations in data collected from the same subject across different time periods.
Table 9: Cross-subject single-session 10 fold cross-validation on MDD with different initial classifiers.
@@ 4.4 Experimental Results on Cross-Subject Cross-Session
AUROC 82.15 82.15 82.15 82.15 80.40 80.40
NPV 82.38 82.38 82.38 82.38 78.86 78.86
EO Accuracy 74.96 74.96 78.63 74.96 74.96 74.96
Sensitivity 73.98 73.98 77.68 73.98 73.98 73.98
Specificity 75.79 75.79 79.24 75.79 75.79 75.79
Precision 73.97 73.97 77.67 73.97 73.97 73.97
@@
AUROC 74.89 74.89 78.46 74.89 74.89 74.89
NPV 75.80 75.80 79.25 75.80 75.80 75.80
Table 10: Performance comparison on MDD using cross-subject 10-fold cross-validation. EC: closed-eye; EO: open-eye. The second-best traditional machine learning method result is underlined. Models reproduced by the authors are indicated with an asterisk (*).
@@
      M3D       82.72/11.35       78.63/13.79
For the SEED-IV database, the performance across various initial classifiers is relatively consistent, with the highest accuracy reaching 64.9% when DT is employed as the initial classifier. Table 7 reports the performance comparison results against the existing machine learning and deep learning methods. Compared to the best results reported in various literature on machine learning methods, the proposed M3D shows a performance improvement of 0.46%. Furthermore, it performs comparably or even outperforms some deep learning methods, including MetaEmotionNet [55] and DAN [14]. The superior cross-session validation performance observed in SEED-IV is largely driven by its smaller per-session sample size, higher emotional class complexity, and more diverse multimodal stimuli, which collectively benefit from cross-session aggregation. Meanwhile, the SEED database’s more stable per-session paradigm and stimulus uniformity make single-session validation more effective and reliable.
Table 11: Performance comparison without classifier learning and ensemble learning, with only classifier learning, and with both classifier learning and ensemble learning.
Manifold Feature Transformation Dynamic Distribution Alignment and Classifier Learning Ensemble Learning Accuracy Sensitivity Specificity Precision F1-score AUROC NPV 48.27/14.39 37.70 54.65 33.40 35.42 61.52 59.26 69.51/16.40 62.64 72.69 51.51 56.53 77.16 80.77 71.01/13.42 65.17 73.83 54.58 59.41 78.25 81.46 79.66/09.34 74.37 82.21 66.86 70.41 84.68 86.92 84.57/09.49 76.04 88.78 76.94 76.49 88.36 88.27
For the SEED-V database in the cross-session setting, performance across various initial classifiers remains relatively stable, with the highest accuracy reaching 59.15%, again observed with DT. As shown in Table 8, the proposed M3D model outperforms the best results reported in the literature for traditional machine learning methods by a margin of 11.69%, demonstrating a substantial improvement in generalization capability. Furthermore, M3D delivers performance that is comparable to or even surpasses certain deep learning models, including the state-of-the-art DCORAL method [59].
4.5 Experimental Results on clinical EEG database
To assess the generalizability of the proposed M3D model in real-world clinical applications, we evaluate its performance on the MDD database (n=56). Following previous work [62], we adopt the cross-subject single-session 10-fold cross-validation protocol for fair comparison. The validation results are presented in Table 9, demonstrating consistent performance across various initial classifiers. In particular, the highest accuracy is achieved using DT as the initial classifier, reaching 82.72% in the EC session and 78.63% in the EO session. For comparative analysis, we report performance differences against existing methods in Table 10. The proposed M3D model outperforms the best-performing traditional machine learning approach by 8.51% and 4.42% in the EC and EO sessions, respectively. Among deep learning methods, DCORAL [59] achieves the highest reported accuracy in prior studies, with 70.88±10.53% in the EC session and 74.49±8.17% in the EO session. M3D surpasses these results by 11.84% and 4.14% in the EC and EO sessions, respectively, demonstrating its superior ability to capture the complex patterns associated with MDD. These results underscore the effectiveness of M3D in real-world clinical settings, establishing it as a promising tool for practical clinical applications.
5 Discussion
5.1 Ablation Study
@@
Table 11 presents the results of the ablation study, demonstrating that the proposed M3D model achieves the best performance by effectively integrating all modules. Using only the manifold feature transformation module leads to the lowest performance, with a classification accuracy of 44.73±14.39%. Focusing on dynamic distribution alignment and classifier learning improves the accuracy to 69.51±16.4%. Adding ensemble learning to dynamic distribution alignment and classifier learning further increases the accuracy to 71.01±13.42%, a 1.5% improvement, highlighting the effectiveness of ensemble learning in EEG-based emotion recognition. However, removing the manifold feature transformation module causes a significant accuracy drop from 84.57±9.49% (full model) to 71.01±13.42%, emphasizing the importance of feature transformation before distribution alignment. Removing the ensemble learning module also reduces the accuracy to 79.66±9.34%. Additionally, features in the Grassmann manifold space prove more robust than those in the original feature space, improving distribution alignment. All four ablation models significantly differ from the full model (p $<$ 0.0001). We use an independent t-test for normally distributed groups and the Wilcoxon Rank Sum test for nonnormally distributed ones, with FDR correction applied to enhance result reliability.
Figure 3: Visualization of topographic analysis, histogram distribution of mutual information between EEG patterns and model predictions, and confusion matrices before and after alignment (a) before alignment, (b) after alignment.
5.2 The effect of the parameter
We conduct a series of experiments under the cross-subject single-session setting using the SEED database to investigate the impacts of key parameters (subspace dimension $d$, regularization parameters $\eta$, $\rho$, $\lambda$), the number of iterations $\iota$, and the training sample size on the performance of the M3D model. The results show that the model achieves the best performance when the subspace dimension $d=128$, and it demonstrates good robustness to variations in the regularization parameters within specific ranges. The model’s classification performance fluctuates at the beginning of iterations and stabilizes after $\iota=10$, demonstrating the training advantage of M3D in cross-domain tasks. Using only 3,000 training samples, the model achieves an accuracy of 83.11% with a runtime of 37.4 seconds, outperforming several deep learning methods that use all available training samples. For instance, GECNN [48] achieves an accuracy of 82.46±10.83% and MS-MDA [45] obtains 79.67±8.01%. More detailed descriptions and results are presented in Supplementary Materials Appendix C. And we provide a comprehensive theoretical analysis of the computational complexity of the proposed M3D framework in Supplementary Materials Appendix D.
5.3 Investigating factors affecting model performance
We also conduct a series of experiments to study the factors that affect the performance of the model. First, replacing the Grassmann manifold with the Stiefel manifold in the manifold feature transformation module leads to a decrease in the model’s accuracy to 67.94%, demonstrating the Grassmann manifold’s superiority in extracting EEG emotion features. Second, comparing TCA with Principal Component Analysis (PCA) in manifold feature transformation, TCA shows better performance in terms of accuracy and F1-score, as it can map, reduce dimensions, and maximize the similarity between domains simultaneously. Third, experiments with fixed $\mu$ values in dynamic distribution alignment show that the adaptive $\mu$ significantly improves the model performance, increasing the accuracy by 4.26%. This mechanism mimics the brain’s ability to prioritize different information streams based on contextual discrepancies, representing a more biologically plausible and data-responsive adaptation process. Finally, the evaluation of ensemble learning indicates that both the voting method and the LinkCluE method can enhance the reliability of classification results, with the LinkCluE method outperforming the voting method. More detailed descriptions and results are presented in the Supplementary Materials Appendix E.
5.4 Visualization of Feature Alignment
We employ t-SNE visualizations and EEG topographic maps to intuitively demonstrate the alignment of source and target features before and after adaptation. For t-SNE [63] visualization, we project the source and target features into a two-dimensional space before and after alignment. Distinct colors represent different emotion categories, while different shapes indicate domain origin (source vs. target). The visualization (Supplementary Materials Appendix F, Fig. S13) clearly demonstrates that after alignment, the distributional discrepancy between the source and target domains is significantly reduced, while the separability between different emotion classes improves.
Additionally, we conduct topographic analysis by computing the mutual information between EEG patterns and prediction labels to identify important brain patterns for emotion recognition. The mutual information between the input features of the target domain and the model prediction results is estimated using a non-parametric method [64, 65, 66]. After normalizing the mutual information matrix, we visualize the average mutual information before and after alignment. The results in Fig. 3 show that the differences among the EEG topographic maps of the three emotions increase after alignment. The confusion matrix further supports this finding, demonstrating that the model achieves better classification performance following alignment. Moreover, the analysis reveals that the EEG patterns with higher informativeness for emotion recognition are mainly located in the prefrontal region [67, 68]. More detailed descriptions and results are presented in the Supplementary Materials Appendix F.
6 Conclusion
In this paper, we propose a novel non-deep transfer learning framework, termed M3D, for cross-subject and cross-session EEG-based emotion recognition, as well as clinical mental disorder recognition. The framework employs a manifold-based domain adaptation with dynamic distribution alignment, adaptively balancing marginal and conditional distributions in the Grassmann manifold space. This enhances intrinsic data representation and reduces domain disparity. Additionally, optimized classifier learning with ensemble techniques improves robustness and reliability in handling EEG data variations. Extensive experiments on three benchmark databases validate the framework under two evaluation protocols. Furthermore, we demonstrate the model’s effectiveness on a real clinical MDD dataset. Compared to existing methods, M3D achieves promising results against non-deep approaches and performance comparable to deep learning models, with an average improvement of 6.67% in classification accuracy. These findings highlight the potential of non-deep transfer learning in addressing individual and session variations in affective brain-computer interfaces. On the other hand, in the dynamic distribution alignment module, we did not impose constraints on the adaptive factor $\mu$ during the iteration process, which may lead to suboptimal feature alignment in certain cases. Addressing this limitation in future work could further enhance the stability and generalizability of the framework. And we plan to investigate the adaptability of the M3D framework to more diverse clinical EEG datasets (e.g., epilepsy monitoring or sleep staging) and to evaluate the stability of model performance under varying preprocessing pipelines.
[…]
7 Proof of the computation $\mu$
The a daptive factor $\mu$ is designed to dynamically adjust the emphasis between marginal and conditional distribution adaptation based on the differences between the source and target domains. We use the A-distance measurement [32] as a metric to estimate these differences, which is a well-established measure in the domain adaptation literature. The adaptive factor $\mu$ is computed as
$\mu=1-\dfrac{d_{A}}{d_{A}+\sum^{C}_{c=1}d_{c}},$ (28)
where $d_{A}$ is the marginal distribution difference between the source and target domains and $d_{c}$ is the conditional distribution difference between the class $c$ ($c\in[1,\dots,C]$) from the source and target domains. When $d_{A}$ is large compared to $\sum^{C}_{c=1}d_{c}$ , it means the marginal distribution difference between the source and target domains is significant. In this case, $\mu$ will be close to 0, indicating that more weight should be placed on marginal distribution adaptation. Conversely, when $\sum^{C}_{c=1}d_{c}$ is large relative to $d_{A}$, $\mu$ will be close to 1, suggesting that conditional distribution adaptation is more crucial.
For the marginal distribution difference, the A-distance $d_{A}$ provides a quantification of the marginal distribution difference between the source domain $T_{S}$ and the target domain $T_{T}$, and is defined as
where $\epsilon\left(h\right)$ denotes the hinge loss, which is derived from a binary classifier trained specifically to distinguish between samples from the source and target domains. A smaller A-distance indicates a closer marginal distribution.
For the conditional distribution difference, based on the provided label information for source data and the estimated pseudo-label information of target data, we conduct the A-distance measurement within each $c$ class as
$d_{c}=d_{A}(T^{{c}}_{S},T^{(c)}_{T}),$ (30)
Here, $T_{S}^{(c)}$ and $T_{T}^{(c)}$ represent the subsets of data corresponding to class $c$ from the source and target domains, respectively. By aggregating these class-specific A-distances ($\sum^{C}_{c=1}d_{c}$) and comparing them with the marginal A-distance $d_{A}$, we can determine the relative importance of marginal and conditional distribution adaptation.
8 Statistically analyze EEG signal differences across subjects
To thoroughly investigate EEG signal differences across subjects in the SEED and SEED-IV datasets, we perform a comprehensive statistical analysis on the input EEG features, which includes both qualitative visualizations and quantitative comparisons. Specifically, our analysis includes three main parts: (1) low-dimensional feature visualization using t-Distributed Stochastic Neighbor Embedding (t-SNE) to reveal the overall data distribution and inter-subject separability; (2) box plot analysis to display the statistical distribution of EEG features for each subject across different emotional states; and (3) statistical difference matrices constructed through rigorous hypothesis testing to quantify the inter-subject variability under the same emotion condition.
Firstly, we apply t-SNE to visualize the distribution of the input EEG features in a two-dimensional space. The results (Fig. 4 and 5) clearly show distinct clustering patterns for different emotional states and reflect the variability of feature distributions across subjects. This visualization offers an intuitive understanding of how subject-specific factors impact feature separability and emotional clustering.
Figure 4: A t-SNE visualization of the input EEG features in SEED database. The blue, green, and red colors indicate negative, neutral, and positive emotions.
Figure 5: A t-SNE visualization of the input EEG in SEED-IV database. The blue, green, red and purple colors indicate negative, neutral, positive and fear emotions.
Secondly, we perform feature dimensionality reduction followed by box plot analysis for each subject and emotion category (Fig. 6-9). The box plots provide detailed statistical information, including medians, interquartile ranges, and outliers, and the analysis results reflect the extent of variation in input EEG features among individuals. We observe that, even within the same emotional condition, the input EEG feature distributions vary significantly across subjects, suggesting that individual differences play a substantial role in neural responses to emotional stimuli, and further, an enlarged intra-subject variance is observed under cross-session settings, indicating temporal fluctuations in individual neural patterns.
Figure 6: Box plots of the input EEG feature distribution across 15 subjects and 3 emotions (negative, neutral, positive) in a single session of SEED database.
Figure 7: Box plots of the input EEG feature distribution across 15 subjects and 3 emotions (negative, neutral, positive) in the cross session of SEED database.
Figure 8: Box plots of the input EEG feature distribution across 15 individuals and 4 emotions (negative, neutral, positive, fear) in a single session of SEED-IV database.
Figure 9: Box plots of the input EEG feature distribution across 15 individuals and 4 emotions (negative, neutral, positive, fear) in the cross session of SEED-IV database.
Thirdly, to systematically quantify the degree of inter-subject difference, we perform statistical difference matrices using independent sample t-tests (or Wilcoxon rank-sum tests when normality assumptions are violated), with all p-values adjusted using False Discovery Rate (FDR) correction. The resulting matrices (Fig. 10-13) indicate statistically significant differences ($p<0.05$) between subjects under the same emotion condition, underscoring the presence of subject-specific EEG patterns. These inter-individual differences may stem from a range of factors, including variations in physiological structure, psychological characteristics, and individual cognitive processing mechanisms.
Figure 10: Statistical difference matrix across subjects in a single session of SEED database, with independent samples t-test and FDR correction (a) Negative, (b) Neutral, (c) Positive.
Figure 11: Statistical difference matrix across subjects and across session of SEED database, with independent samples t-test and FDR correction (a) Negative, (b) Neutral, (c) Positive.
Figure 12: Statistical difference matrix across subjects in a single session of SEED-IV database, with independent samples t-test and FDR correction (a) Negative, (b) Neutral, (c) Positive, (d) Fear.
Figure 13: Statistical difference matrix across subjects and across session of SEED-IV database, with independent samples t-test and FDR correction (a) Negative, (b) Neutral, (c) Positive, (d) Fear.
In summary, our analysis based on three parts (low-dimensional visualizations using t-SNE, distributional comparisons through box plots, and statistical testing via difference matrices) demonstrates that EEG signals exhibit strong individual variability across subjects and emotional conditions, both visually and statistically. These findings highlight the complexity of cross-subject EEG emotion recognition and emphasize the need for robust adaptation mechanisms such as our proposed M3D framework.
9 The effect of the parameter
We conduct supplementary experiments under the cross-subject single-session setting using the SEED database to discuss the impacts of key parameters (subspace dimension $d$, regularization parameters $\eta$, $\rho$, $\lambda$), the number of iterations $\iota$, and the training sample size on the performance of the M3D model. The subspace dimension $d$ is varied across $[10,24,36,48,64,128,256,310]$, while the regularization parameters $\eta$, $\rho$, and $\lambda$ are explored at multiple values within the ranges of $[0.001-100]$, $[0.001-50]$, and $[0.001-100]$, respectively. As shown in Fig. 14, the experimental results demonstrate that the M3D model achieves optimal performance across multiple evaluation metrics, when $d$ is set to 128, $\eta$ to 0.1, $\rho$ to 1, and $\lambda$ to 0.4. Additionally, the model maintains stable performance across a broad range of parameter values, particularly when $\eta\in[0.05,1]$, $\rho\in[0.1,1]$, and $\lambda\in[0.1,1]$. These findings confirm the strong parameter adaptability of the proposed M3D model and offer practical guidance for hyperparameter tuning in real-world applications.
Additionally, to examine the classifier learning process, we also investigate how sensitive the model is to changes in the number of iterations ($\iota$) in the learning process. Here, we vary $\iota$ value in a range of $[2,5,10,20,30,50]$. The corresponding learning results, expressed as accuracy rates, are 80.39%, 83.81%, 84.57%, 83.84%, 83.34% and 82.55% respectively. It shows the model’s classification performance experiences fluctuations during the initial few iterations but achieves stability beyond $\iota=10$ iterations. This demonstrates the training advantage of M3D in cross-domain tasks.
We also explore the impact of the training sample size on the model’s performance, as well as their computation time. We train the model using training set samples of different scales, with the number of training samples ranging from 1,000 to 20,000. The experimental results in Fig. 15 and Table 12 show that our method achieves an accuracy of 83.11% with only 3,000 training samples, and the runtime is only 37.4 seconds. Notably, this outperforms several deep learning methods that use all available training samples. For instance, GECNN [48] achieves an accuracy of 82.46±10.83%, MS-MDA [45] obtains 79.67±8.01%, and R2G-STLT [46] reaches 77.96±6.38%.
Figure 14: Parameter sensitivity and convergence analysis of M3D (a) subspace dimension $d$, (b) regularization parameters $\eta$, (c) regularization parameters $\rho$, (d) regularization parameters $\lambda$.
Table 12: Running time (s) of the proposed M3D under different sample sizes.
Sample size 1000 2000 3000 4000 5000 6000 7000 10000 15000 20000
Runnning time (s) 19.32 27.63 37.40 53.35 81.81 88.51 100.95 263.17 449.41 1205.00
Table 13: Performance comparison using different modules.
Manifold Stiefel manifold Grassmann manifold
Accuracy 67.94 84.57
Sensitivity 63.30 76.04
Specificity 70.37 88.78
Precision 52.77 76.94
F1-Score 57.55 76.49
AUROC 75.95 88.36
NPV 78.57 88.27
Dimensionality Reduction PCA TCA
Accuracy 56.81 84.57
Sensitivity 77.73 76.04
Specificity 50.91 88.78
Precision 30.86 76.94
F1-Score 44.18 76.49
AUROC 78.28 88.36
NPV 89.03 88.27